# The general Brannan coefficient conjecture and Watson’s lemma

### v1

@@ 4 Proof of positivity for $n=5$
Let $w_{n}(\alpha,\beta,x)$ be the scaled hypergeometric function defined by Eq. 12, $A_{1,2}(\alpha,\phi)$ be given by Eqs. 52 and 53, and $m_{0}$ and $m_{\infty}$ be the infima given by Eqs. 76 and 77. Then from Eq. 28, the scaling Eq. 55 together with Euler’s reflection identity Eq. 8, and the decomposition Eqs. 54 and 56 combined with the bounds Eqs. 78 and 79, we deduce the following.
@@
For $n=1,3,5,\dots$, $\alpha,\beta\in(0,1]$ and $\phi\in[\phi_{0},\pi)$
@@
(85) $Q_{n}(\alpha,\beta,\phi)=\frac{\pi\,(\alpha+\beta)\sin(\pi\,\alpha)n!(\pi-\phi)^{2}}{(1-\alpha)\Gamma(\beta)\Gamma(n-\alpha)n^{1+\alpha-\beta}}.$
@@
In addition, $P_{n}(\alpha,\beta,\pi)$ is given by Eq. 84 with $A_{1}(\alpha,\pi)$ and $A_{2}(\alpha,\beta,\pi)$ being given by Eqs. 62 and 65, and incidentally from this (for $n=5$) we find numerically that
@@
See Figs. 3 and 4 for surface plots of $P_{5}(\alpha,\beta,\phi_{0})$ (the latter depicting the surface close to the minimum). That the (positive) infimum Eq. 91 is relatively small is a consequence of our choice of $\phi_{0}=0.061$ being as small as possible. We chose the smallest such value (to three digits) that would still result in Eq. 91 being positive. We see from Eq. 53 that $A_{2}(\alpha,\beta,\phi)\to-\infty$ as $\phi\to 0^{+}$. In addition, the infima $m_{0}$ and $m_{\infty}$, as given by Eqs. 59, 72, 73, 74, and 75, also depend on $\phi_{0}$, and they too approach -$\infty$ as $\phi_{0}\to 0^{+}$. Hence from Eq. 84 $P_{5}(\alpha,\beta,\phi)\to$-$\infty$ in this limit (noting from Eq. 52 that $A_{1}(\alpha,\beta,\phi)$ remains bounded).
To numerically obtain the infimum of $P_{5}(\alpha,\beta,\phi)$ given in Eq. 84 we carried out a 2-D $\alpha,\beta$ grid search in Maple (using high-precision arithmetic), storing at each step the smallest current value of $P_{5}(\alpha,\beta,\phi)$. We fixed $\phi_{0}=0.061$ and $n=5$, and used the numerical values of $m_{0}$ and $m_{\infty}$ from Eqs. 76 and 77. The search sampled $\alpha$ and $\beta$ over $60^{2}$ point pairs on a grid biased toward 0 and $1$ (including boundary values).
For each grid pair $(\alpha,\beta)$ we minimized over $\phi\in[\phi_{0},\pi]$ by evaluating $P_{5}(\alpha,\beta,\phi)$ only at the endpoints $\phi=\phi_{0}$ and $\phi=\pi$, together with any interior stationary points in $(\phi_{0},\pi)$ where $\partial P_{5}(\alpha,\beta,\phi)/\partial\phi=0$. These stationary points were computed from exact closed-form formulas for the derivatives, using fsolve with sign-change bracketing. Since the cases $\alpha,\beta\in\{0,1\}$ and $\phi=\pi$ are removable singularities, near-edge values were snapped to $\alpha,\beta=0$ or $1$, and $\phi=\pi$, with both $P_{5}(\alpha,\beta,\phi)$ and $\partial P_{5}(\alpha,\beta,\phi)/\partial\phi$ being evaluated via the corresponding exact limiting formulas on edges and corners (see Eqs. 62, 65, 86, 87, 88, 89, and 90). Over the resulting candidate set in $\phi$, the smallest value recorded on the $(\alpha,\beta)$ grid yielded Eq. 91, attained as stated at $(\alpha,\beta,\phi)=(0,1,\phi_{0})$. The exact value shown comes from Eq. 87.
@@ 5 Proof for $n=7,9,11,13,\ldots$
The purpose of this section is to prove the following, which in conjunction with Eqs. 83, 91, and 95, proves that $w_{n}(\alpha,\beta,-1)-|w_{n}(\alpha,\beta,e^{i\phi})|\geq 0$ for $\phi\in[\phi_{0},\pi]$ and $n=7,9,11,\ldots$.
@@ 5.1 Case 1: $c_{2}(\alpha,\beta)\geq 0\;(\alpha\geq\beta)$
Since $A_{1}(\alpha,\phi)$ and $A_{2}(\alpha,\beta,\phi)$ are smooth in $(\alpha,\beta,\phi)$ and from Eqs. 60 and 63
@@ 5.2 Case 2: $c_{2}(\alpha,\beta)<0\;(\alpha<\beta)$
Therefore, in conjunction with Eqs. 99 and 100, it is evident that
@@
Thus, in summary, to establish $\partial J/\partial n\geq 0$ for $n\geq 5$ it is sufficient from Eqs. 101 and 103 to check that $F_{1}(\alpha,\beta,\phi)\geq 0$ for $0\leq\beta\leq\alpha\leq 1$, and $F_{2}(\alpha,\beta,\phi)\geq 0$ for $0\leq\alpha<\beta\leq 1$. The following then completes the proof of Theorem 5.1.
@@
The computation of these values was performed in a similar manner to the method used for Proposition 4.2, that is, sampling over a fine $\alpha,\beta$ grid, and at $\phi$ boundary points ($\phi_{0}$ and $\pi$) and at any interior critical points. The infima were then independently confirmed using the Maple Optimization[Minimize] package. The grid and optimization code can be found in [Dunster:2026:BWM].
@@
where $A_{1}(\alpha,\pi)$, $A_{2}(\alpha,\beta,\pi)$ and $c_{2}(\alpha,\beta)$ are given by Eqs. 62, 65, and 94.
Incidentally, in comparison to Eqs. 105 and 106, from Eqs. 111 and 112 we find
@@ 6 Extension to $\phi\in[0,\phi_{0})$ ($\theta\in(\pi-\phi_{0},\pi]$)
We finish by noting the following simple result, which confirms without using Kummer functions that Eq. 5 holds at $\theta=\pi$ ($\phi=0$) provided $\beta\geq\alpha$.
@@
Let $0<\alpha\leq\beta<1$ and let $1\leq n\in\mathbb{N}$ be odd. Then
@@
Thus for $\alpha\in(0,1)$ the $k=0$ and the $k$-odd coefficients in Eq. 122 are positive. Therefore, using Eq. 122 with $\omega=\pm 1$ and the fact that $n$ is assumed to be odd, we obtain for $0<\alpha\leq\beta<1$

### v2

@@ 4 Proof of positivity for $n=5$
Let $w_{n}(\alpha,\beta,x)$ be the scaled hypergeometric function defined by Eq. 10, $A_{1,2}(\alpha,\phi)$ be given by Eqs. 53 and 54, and $m_{0}$ and $m_{\infty}$ be the infima given by Eqs. 77 and 78. Then from Eq. 29, the scaling Eq. 56 together with Euler’s reflection identity Eq. 8, and the decomposition Eqs. 55 and 57 combined with the bounds Eqs. 79 and 80, we deduce the following.
@@
For $n=1,3,5,\dots$, $\alpha,\beta\in(0,1]$ and $\phi\in[\phi_{0},\pi]$
@@
(86) $Q_{n}(\alpha,\beta,\phi)=\frac{\pi\,(\alpha+\beta)\sin(\pi\alpha)\,\Gamma(1+\alpha)\,(\pi-\phi)^{2}}{(1-\alpha)\Gamma(\beta)\,n^{1+\alpha-\beta}}\geq 0.$
@@
In addition, $P_{n}(\alpha,\beta,\pi)$ is given by Eq. 85 with $A_{1}(\alpha,\pi)$ and $A_{2}(\alpha,\beta,\pi)$ being given by Eqs. 63 and 66, and incidentally from this (for $n=5$) we find numerically that
@@
See Figs. 3 and 4 for surface plots of $P_{5}(\alpha,\beta,\phi_{0})$ (the latter depicting the surface close to the minimum). That the (positive) infimum Eq. 92 is relatively small is a consequence of our choice of $\phi_{0}=0.061$ being as small as possible. We chose the smallest such value (to three digits) that would still result in Eq. 92 being positive. We see from Eq. 54 that $A_{2}(\alpha,\beta,\phi)\to-\infty$ as $\phi\to 0^{+}$. In addition, the infima $m_{0}$ and $m_{\infty}$, as given by Eqs. 60, 73, 74, 75, and 76, also depend on $\phi_{0}$, and they too approach -$\infty$ as $\phi_{0}\to 0^{+}$. Hence from Eq. 85 $P_{5}(\alpha,\beta,\phi)\to$-$\infty$ in this limit (noting from Eq. 53 that $A_{1}(\alpha,\beta,\phi)$ remains bounded).
To numerically obtain the infimum of $P_{5}(\alpha,\beta,\phi)$ given in Eq. 85 we carried out a 2-D $\alpha,\beta$ grid search in Maple (using high-precision arithmetic), storing at each step the smallest current value of $P_{5}(\alpha,\beta,\phi)$. We fixed $\phi_{0}=0.061$ and $n=5$, and used the numerical values of $m_{0}$ and $m_{\infty}$ from Eqs. 77 and 78. The search sampled $\alpha$ and $\beta$ over $60^{2}$ point pairs on a grid biased toward 0 and $1$ (including boundary values).
For each grid pair $(\alpha,\beta)$ we minimized over $\phi\in[\phi_{0},\pi]$ by evaluating $P_{5}(\alpha,\beta,\phi)$ only at the endpoints $\phi=\phi_{0}$ and $\phi=\pi$, together with any interior stationary points in $(\phi_{0},\pi)$ where $\partial P_{5}(\alpha,\beta,\phi)/\partial\phi=0$. These stationary points were computed from exact closed-form formulas for the derivatives, using fsolve with sign-change bracketing. Since the cases $\alpha,\beta\in\{0,1\}$ and $\phi=\pi$ are removable singularities, near-edge values were snapped to $\alpha,\beta=0$ or $1$, and $\phi=\pi$, with both $P_{5}(\alpha,\beta,\phi)$ and $\partial P_{5}(\alpha,\beta,\phi)/\partial\phi$ being evaluated via the corresponding exact limiting formulas on edges and corners (see Eqs. 63, 66, 87, 88, 89, 90, and 91). Over the resulting candidate set in $\phi$, the smallest value recorded on the $(\alpha,\beta)$ grid yielded Eq. 92, attained as stated at $(\alpha,\beta,\phi)=(0,1,\phi_{0})$. The exact value shown comes from Eq. 88.
@@ 5 Proof for $n=7,9,11,13,\ldots$
The purpose of this section is to prove the following, which in conjunction with Eqs. 84, 92, and 96, proves that $w_{n}(\alpha,\beta,-1)-|w_{n}(\alpha,\beta,e^{i\phi})|\geq 0$ for $\phi\in[\phi_{0},\pi]$ and $n=7,9,11,\ldots$.
@@ 5.1 Case 1: $c_{2}(\alpha,\beta)\geq 0\;(\alpha\geq\beta)$
Since $A_{1}(\alpha,\phi)$ and $A_{2}(\alpha,\beta,\phi)$ are smooth in $(\alpha,\beta,\phi)$ and from Eqs. 61 and 64
@@ 5.2 Case 2: $c_{2}(\alpha,\beta)<0\;(\alpha<\beta)$
Therefore, in conjunction with Eqs. 100 and 101, it is evident that
@@
Thus, in summary, to establish $\partial J/\partial n\geq 0$ for $n\geq 5$ it is sufficient from Eqs. 102 and 104 to check that $F_{1}(\alpha,\beta,\phi)\geq 0$ for $0\leq\beta\leq\alpha\leq 1$, and $F_{2}(\alpha,\beta,\phi)\geq 0$ for $0\leq\alpha<\beta\leq 1$. The following then completes the proof of Theorem 5.1.
@@
The computation of these values was performed in a similar manner to the method used for Proposition 4.2, that is, sampling over a fine $\alpha,\beta$ grid, and $\phi$ only at boundary points ($\phi_{0}$ and $\pi$) and at any interior critical points. The infima were then independently confirmed using the Maple Optimization[Minimize] package. The grid and optimization code can be found in [Dunster:2026:BWM].
@@
where $A_{1}(\alpha,\pi)$, $A_{2}(\alpha,\beta,\pi)$ and $c_{2}(\alpha,\beta)$ are given by Eqs. 63, 66, and 95.
Incidentally, in comparison to Eqs. 106 and 107, from Eqs. 112 and 113 we find
@@ 6 Extension to $\phi\in[0,\phi_{0})$ ($\theta\in(\pi-\phi_{0},\pi]$)
We finish by noting the following simple result, which confirms without using Kummer functions that Eq. 5 holds ($n$ even or odd) at $\theta=\pi$ ($\phi=0$), provided $\beta\geq\alpha$.
@@
Let $0<\alpha\leq\beta<1$ and let $1\leq n\in\mathbb{N}$. Then
@@
Thus for $\alpha\in(0,1)$ the $k=0$ and the $k$-odd coefficients in Eq. 123 are positive. Therefore, using Eq. 123 with $\omega=\pm 1$, we obtain for $0<\alpha\leq\beta<1$